A food safety guideline is that the mercury in fish should be below 1 part per million​ (ppm). Listed below are the amounts of mercury​ (ppm) found in tuna sushi sampled at different stores in a major city. Construct a

98​%

confidence interval estimate of the mean amount of mercury in the population. Does it appear that there is too much mercury in tuna​ sushi?

0.52 

0.80 

0.11 

0.99  

1.26 

0.55 

0.93

What is the confidence interval estimate of the population mean

muμ​?
Answer #1

Confidence interval for Population mean is given as below:

Confidence interval = Xbar ± t*S/sqrt(n)

From given data, we have

Xbar = 0.737142857

S = 0.37725828

n = 7

df = n – 1 = 6

Confidence level = 98%

Critical t value = 3.1427

(by using t-table)

Confidence interval = Xbar ± t*S/sqrt(n)

Confidence interval = 0.737142857 ± 3.1427*0.37725828/sqrt(7)

Confidence interval = 0.737142857 ± 0.4481

Lower limit = 0.737142857 - 0.4481 = 0.2890

Upper limit = 0.737142857 + 0.4481 = 1.1853

Confidence interval = (0.2890, 1.1853)

It does not appear that there is too much mercury in tuna​ sushi, because the above interval contains the value one.